The estimated lifetime of the Sun on the main sequence is ~ 1010 years. Calculate the main sequence lifetime of stars of masses (i) 100 MΘ and (ii) 5 MΘ.
1
Expert's answer
2019-04-02T10:41:01-0400

An expression for the main sequence life-time is:


"\frac{t_{MS}}{t_{Sun}}={\frac{M_{MS}}{M_{Sun}}}^{-2.5}"

So, is we know relative star mass, life-time is estimated to be:


"t_{MS}=t_{Sun}{\frac{M_{MS}}{M_{Sun}}}^{-2.5}"

i)

"t_{MS}=t_{Sun}{\frac{M_{MS}}{M_{Sun}}}^{-2.5}=10^{10}*{100}^{-2.5}=10^5 years"

ii)

"t_{MS}=t_{Sun}{\frac{M_{MS}}{M_{Sun}}}^{-2.5}=10^{10}*{5}^{-2.5}=1.8*10^8 years"